OpenStudy (anonymous):

If 24 is increased to 30, the percent of increase is a0%.

OpenStudy (anonymous):

increase=30-24=6 percent increase=6/24x100% = 25%

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

my plzre

OpenStudy (anonymous):

:)